You are bearish on Telecom and decide to sell short 100 shares at the current market price of $50 per share. How much in cash or securities must you put into your brokerage account if the brokers initial margin requirement is 50% of the value of the short position? Enter your answer without a dollar sign Your Answer: Answer Question 5 (1 point) continued from the previous question) How high can the price of the stock go before you get a margin call if the maintenance margin is 25% of the value of the short position? Round your answer to two decimal places and enter it without dollar sign. Your Answer: Answer
Answer #1

Value of short position = 100 * 50 = 5000

If broker initial margin requirements is 50% then amount required in cash or securities = 5000*50% = 2500

Maintenance margin = 25%

Margin currently in account = 2500

Calculating how high the price of stock can go before you get a margin call = (2500*100/25)/100 shares = 10000/100 = $100 per share